What is the value of ? for this aqueous reaction at 298 K?

A+B↽−−⇀C+DΔ?°=29.94 kJ/mol A + B ↽ − − ⇀ C + D Δ G ° = 29.94 kJ / mol
Answer #1

T = 298 K

ΔGo = 29.94 KJ/mol

ΔGo = 29940 J/mol

use:

ΔGo = -R*T*ln Kc

29940 = - 8.314*298.0* ln(Kc)

ln Kc = -12.0844

Kc = 5.647*10^-6

Answer: 5.647*10^-6